OpenStudy (anonymous):

Explain why (–4x)0 = 1, but –4x0 = –4.

OpenStudy (anonymous):

I think you mean (-4x)^0 =1 because a^0 =1 , wheras -4x^0 =-4*1=-4

OpenStudy (anonymous):

I think you mean (-4x)^0 =1 because a^0 =1 , wheras -4x^0 =-4*1=-4